OpenStudy (anonymous):

What is the domain of f(x) = sqrt x - 7? Help!!

OpenStudy (anonymous):

all values of x that will make the expression inside the square root sign nonnegative. Is it \[f(x)=\sqrt x-7\]or \[f(x)=\sqrt{x-7}\]

OpenStudy (anonymous):

its the second one

OpenStudy (anonymous):

ok... :D now let see. from what i have said previously, we must have \[x-7\ge0\]then\[x \ge 7\]

OpenStudy (anonymous):

that's what I got lol